Why do we need a mathematical equation to tell us how many heterozygous carriers there are in a population?
Answer #1

According to biological concept of hardy weinberg equilibrium,

p + q = 1

Where, p is the frequency of dominant allele and q is that of recessive allele. By using this simple equation, we can tell the allelic (p and q) as well as genotypic (pp and qq) frequency of both the alleles and genotypes respectively.

But what about heterozygotes, which carry both the alleles and have a genotype of pq....? So, for this we expand the above by squaring it on both the sides and then wd can find out the genotypic frequency of heterzygotes.

Squaring both sides, we get a mathematical equation :

(p + q)2 = 12

p2 + 2pq + q2 = 1

2pq = 1 - p2 - q2

2pq is the frequency of heterozygotes.

If you find the answer useful then please hit the like button.